OpenStudy (mathmath333):

A man buys milk at Rs. 8.5 per litre and dilutes it with water. He sells the mixture at the same rate and thus gains 11.11%. Find the quantity of water mixed by him in every litre of the milk

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{a.0.111\\ b.0.909\\ c.0.1\\ d.0.125}\end{align}\)

ganeshie8 (ganeshie8):

what do you think ?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\frac{n_1}{n_2}=\frac{A_2-A_w}{A_w-A_1}\\\text{i dont know how to use formula}}\end{align}\)

ganeshie8 (ganeshie8):

save the formulas, this is a simple problem and should be solvable easily..

ganeshie8 (ganeshie8):

since he is mixing water, he must be selling less than 1L at 8.5, yes ?

ganeshie8 (ganeshie8):

since he is mixing water, he must be selling less than 1L milk at 8.5, yes ?

OpenStudy (mathmath333):

yes

ganeshie8 (ganeshie8):

suppose he puts `x` liter pure milk and `y` liter water in every 1 liter of mixture that he sells

ganeshie8 (ganeshie8):

since he is making a profit of 11.11% for every 1L mixture he sells : x*1,1111 = 1

ganeshie8 (ganeshie8):

x = 1/1.1111 = 0.90009 L

ganeshie8 (ganeshie8):

0.9 L of milk of every Liter of mixture

ganeshie8 (ganeshie8):

so how much is the water in every liter of mixture ?

ganeshie8 (ganeshie8):

1 - 0.9 = 0.1 L of water in every 1L of mixture

OpenStudy (mathmath333):

but book is giving 0.111

ganeshie8 (ganeshie8):

books answer is wrong

ganeshie8 (ganeshie8):

0.111 L is the amount of water `for` every 1 liter of milk

ganeshie8 (ganeshie8):

however its not the amount of water `in` every 1 liter of mixture that he sells

ganeshie8 (ganeshie8):

you will get 0.111 L by solving below : say he mixes `x` L of water for ever liter of milk he gets, then for 11.11% for profit : (1+x)8.5 = 8.5 + 11.11% of 8.5 solve you get x = 0.111 http://www.wolframalpha.com/input/?i=solve+%281%2Bx%298.5+%3D+8.5+%2B+0.1111*8.5

ganeshie8 (ganeshie8):

but thats not what was asked in the question

OpenStudy (mathmath333):

ok thanks